OpenStudy (anonymous):

I don,t get integral[sin(x^2)]dx

OpenStudy (turingtest):

Neither do I. It seems to be some kind of answer that cannot be represented in terms of simple functions.

OpenStudy (asnaseer):

look here and click "show steps" to see how it is done: http://www2.wolframalpha.com/input/?i=integrate+sin%5E2%28x%29

OpenStudy (turingtest):

that is a different integral asnaseer

OpenStudy (turingtest):

what do you mean Yahya?\[\int\sin(x^2)dx\]or\[\int\sin^2(x)dx\]???

OpenStudy (asnaseer):

oh - the other form is shown here: http://www2.wolframalpha.com/input/?i=integrate+sin%28x%5E2%29 but it doesn't show the steps of how to get the answer.

OpenStudy (turingtest):

Yes, and I'm unfamiliar with the "Fresnel integral"

OpenStudy (asnaseer):

Fresnel integrals are described here (if you are interested in learning about them): http://itcanbeshown.com/integrals/Fresnel%20Integrals/fresnel_integrals.pdf